OpenStudy (anonymous):

find all vertical asymptotes of the function f(x) = x^2+9x+20/x^2-2x-24

OpenStudy (anonymous):

Vertical Asymptotes occur when the bottom equals 0 so find an x value that would make the denominator equal to 0

OpenStudy (anonymous):

Do you understand? :)

OpenStudy (anonymous):

I don't quite understand how to do it..

OpenStudy (anonymous):

Should I be simplifying it to its foiled form?

OpenStudy (anonymous):

yeah set the bottom to zero \[x ^{2} - 2x - 24 = 0\]

OpenStudy (anonymous):

Now split them up \[(x -6) (x-4) = 0\]

OpenStudy (anonymous):

oops sorry it was (x-6)(x+4) =0

OpenStudy (anonymous):

now set both equal to zero x-6=0 x+4=0

OpenStudy (anonymous):

now solve for x on both and then that should give you the answers :)

OpenStudy (anonymous):

Would I have to do anything with the numerator?

OpenStudy (anonymous):

Nope! The vertical asymtotes occur when the denominator equals zero so your answer would be x=6 and x=-4

OpenStudy (anonymous):

and you can check your work by plugin the Xs back into the denominator and if they equal to zero then they're right :)

OpenStudy (anonymous):

Okay! I understand it much better now, thanks! (:

OpenStudy (anonymous):

Okay, if you need any more help I'll be glad to :)